The manager of a local motel asked you to figure out his cost structure to project his future operations. According to his records, the motel recorded the total costs of $72,000 by selling 7,500 rooms during the busiest month of the past year. During the slowest month, it sold 4,500 rooms at the total costs of $48,000. Based on the given information, please answer the following questions.

Find out the motel’s fixed cost amount per month

Find out the motel’s variable cost per room sold?   
Answer #1
Total Costs Number of Rooms Sold
At High Point $     72,000 7500
At Low Point $     48,000 4500
Change $     24,000 3000

Working:

Variable cost per room sold = $24000/3000 = $8 per room

Fixed cost per month (at the high point) = $72000 - (7500 x $8) = $72000 - $60000 = $12000 per month

Fixed cost per month (at the low point) = $48000 - (4500 x $8) = $48000 - $36000 = $12000 per month